# Bayesian Multifractal Image Segmentation

### v1

@@ 3 Towards Multifractal-based Image Segmentation in the Fourier Domain
Model (2.3) provides an accurate estimation of $c_{2}$ (recall that $-c_{2}={\theta}_{1}$). However, it is not designed for images with several multifractal textures. Here, we propose a new statistical model in the Fourier domain for modeling heterogeneous multifractal images. The limitation of (2.3) for computing $c_{2}$ is its reliance on a regular spatial grid. This prohibits its use for scenarios involving heterogeneous image regions with irregular shapes having different multifractal properties. We therefore develop a spectral likelihood that can handle irregular lattices [35, 36], resulting in a more flexible model.
3.1 Likelihood on Irregular Lattices
Consider an image F that is composed of zones with different multifractal spectra. The log-leaders of F can be assumed to be a mixture of $K$ classes denoted as $\mathcal{C}_{k}$ with different multifractal parameters $\boldsymbol{\theta}_{k}\triangleq[{\theta}_{1}^{k},{\theta}_{2}^{k}]^{T}$, $k=1,...,K$. For segmentation purposes, let $z_{j,n}$ be the label at position $n$ and scale $j$ that maps the observation ${\ell}_{j,n}$ to the class $\mathcal{C}_{k}$, let $\textbf{z}_{j}=[z_{j,1},...,z_{j,N^{2}_{j}}]^{T}$ denotes the label vector and z=[${\textbf{z}}=[\textbf{z}^{T}_{j_{1}},...,\textbf{z}^{T}_{j_{2}}]^{T}$. Let ${I}_{j,k}=\{{n}\in{N}^{2}_{j}\,|{z}_{j,n}=k\}$ be the subset of sites in $\textbf{z}_{j}$ with label $k$ at scale $j$, and $\mathcal{T}_{j,\mathbf{n},k}$ its corresponding indicator function,
@@ 3.1 Likelihood on Irregular Lattices
where $p(\ell_{j,\mathbf{n}\in I_{j,k}}|\boldsymbol{\theta}^{T}_{K})$ is the likelihood of a stationary zero mean Gaussian process with covariance (8) observed at locations $\mathbf{n}\in I_{j,k}$, and $\boldsymbol{\Theta}=[\boldsymbol{\theta}^{T}_{1},...,\boldsymbol{\theta}^{T}_{K}]^{T}$ contains the whole set of multifractal parameters. Just like (7), (15) is problematic for direct use in practice. In addition, obtaining an approximation similar to (2.3) for (15) is less straightforward.
@@ 3.1.1 Debiased Fourier Coefficients
Having different multifractality regions implies rewriting (2.3) using the coefficients that share the same multifractality, yielding irregular sampling in the Fourier domain which requires to adjust for bias [35, 36]. The log-leader Fourier coefficients are redefined for the $k$-th class label in terms of the irregular lattice $\mathcal{T}_{j,{n},k}$ as
@@
where $\xi_{j,k}=1/\big{(}{\sum_{\mathbf{n}}\mathcal{T}^{1/2}_{j,\mathbf{n},k}}\big{)}$ and $t_{j,k}=\xi^{-1}_{j,k}{\sum_{\mathbf{n}}\mathcal{T}_{j,\mathbf{n},k}\boldsymbol{\ell}_{j,\mathbf{n}}}$ are scaling constants considering the bias introduced by the irregular lattice [35, 36].
@@ 3.1.2 Covariance Matrix
where $\mathbf{W}_{j,\textbf{n},k}$ is a weight matrix defined as [35, 36]:
@@
Accordingly, making use of data augmentation as in (2.3), the likelihood (20) can be written in a form that factorizes terms with the parameters ${\theta}^{k}_{1}$ and ${\theta}^{k}_{2}$, as
$p\big{(}{\textbf{x}}_{k},{\boldsymbol{\mu}}_{k}|{\theta}^{k}_{1},{\theta}^{k}_{2},\textbf{z}\big{)}\propto$ $\prod^{S}_{{s}=1}({\theta}^{k}_{2})^{-S}\exp\bigg{(}-\frac{1}{{\theta}^{k}_{2}}\boldsymbol{\mu}_{k}^{H}\mathbf{W}_{s,k}G^{-1}_{2,s}\boldsymbol{\mu}_{k}\bigg{)}$
×$\times({\theta}^{k}_{1})^{-S}\exp\bigg{(}-$ $\frac{1}{{\theta}^{k}_{1}}({\textbf{x}}_{k}-{\boldsymbol{\mu}}_{k})^{H}\mathbf{W}_{s,k}G^{-1}_{1,s}({\textbf{x}}_{k}-\boldsymbol{\mu}_{k})\bigg{)},$ (21)
where $\mathbf{W}_{s,k}\triangleq\mathbf{W}_{j,\textbf{n},k}$, and ${\boldsymbol{\mu}}_{k}=[{{\mu}}_{1,k}^{T},...,{{\mu}}_{S,k}^{T}]^{T}$ is the complex-valued latent vector. The likelihood for all classes is given by
@@ 4 Bayesian Model for Multifractal Segmentation
With the likelihood (22), the multifractal parameters and labels corresponding to the different regularities in the image can be estimated using Bayesian estimators. This approach uses Bayesian inference, where prior distributions are assigned to the unknown model parameters, and these parameters are estimated using the resulting posterior distribution. In this section, the proposed Bayesian model for image segmentation and multifractal parameter estimation is presented.
@@ 4.0.1 Multiscale Label Prior
To account for the multiscale nature of the log-leaders, the corresponding label dependencies within and across scales are modeled via a graph. Specifically, inter-scale connections link neighbors across scales, and the intra-scale connections model the spatial correlation of labels. Figure 1-(a) illustrates the proposed graph structure at a given scale $j$, where the labels to be sampled at position $n$ (white circle) use the values of its neighbors in space and scale (gray dots) to update their values accordingly. Formally, the corresponding multiscale prior for $\textbf{z}_{j}\in\{1,...,K\}^{n}$ at scale $j$ is defined as a modified version of the classical Potts-Markov field [37],
@@
The potential for the spatial (intra-scale) dependencies is defined as
@@
where $\mathcal{V}(n)$ contains the four spatial neighbors of $n$ [38], as depicted in Fig. 1-(a). The potential for the scale (inter-scale) dependencies is defined as
$\Phi_{\beta_{s}}(\textbf{z}_{j},\textbf{z}_{-j})\triangleq\sum_{n=1}^{N^{2}_{j}}$ $\sum_{p\in D(n)}\beta_{s}\,\delta\Big{(}{z}_{j,n}-\textbf{z}_{j+1,p}\Big{)}+$
$\sum_{q\in U(n)}\beta_{s}\,\delta\Big{(}{z}_{j,n}-\textbf{z}_{j-1,q}\Big{)},$ (26)
[…]
@@
which is generally intractable because the number of summation terms grows exponentially with the size of $\textbf{z}_{j}$ [39].
@@ 4.1.1 Granularity Parameters $\beta_{s}$ and $\beta^{j}_{xy}$
4.1.2 Multifractal Parameter Prior
The natural conjugate priors for the parameters $\boldsymbol{\theta}^{k}_{i}$ are inverse-gamma distributions $\mathcal{IG}(\alpha_{i,k},\gamma_{i,k})$ [4, 29], with fixed constants $\alpha_{i,k},\gamma_{i,k}$ $(i=1,2)$ defined, here, for each multifractal region $k=1,...,K$. In practice, these constants are fixed to small values, ensuring that they are close to non-informative Jeffreys priors [4].
4.2 Posterior Distribution
Figure 2: Directed acyclic graph (DAG) of the Bayesian model: the ellipse nodes show the estimated quantities and the square ones, the fixed ones. The FFT transform links the $\boldsymbol{\ell}$ parameters to x and $\boldsymbol{\mu}$.
The posterior distribution for the unknown model parameters can be written in terms of the spectral likelihood (22) and the priors as
@@ 4.2 Posterior Distribution
The Bayesian model is summarized in Fig. 2. Note that given the complexity of the posterior distribution (28), closed-form expressions for the estimators of the unknown parameters $\boldsymbol{\Theta}$, z and $\boldsymbol{\beta}$ cannot be obtained. As an alternative, it is possible to generate samples that are asymptotically distributed according to the target distribution (28) [39]. These samples can be used to approximate Bayesian estimators of the unknown model parameters.
@@ 5 Sampling Procedure
However, z cannot be sampled jointly since its entries are correlated by the Potts prior. The conditional distribution of a single label $z_{j,n}$, $1<j<J$, given all the other labels $\textbf{z}_{\backslash(j,n)}$, is
${z}_{j,{n}}\,|\,\textbf{z}_{\backslash(j,n)},{\textbf{x}},{\boldsymbol{\mu}},\boldsymbol{\Theta},\boldsymbol{\beta}\,\,\propto\\ p({z}_{j,{n}}|\textbf{z}_{\backslash(j,n)},\boldsymbol{\beta})\,\,p\big{(}{\textbf{x}},{\boldsymbol{\mu}}|\boldsymbol{\Theta},\textbf{z}_{\backslash(j,n)}\big{)}$ (29)
@@
$p({z}_{j,{n}}|\textbf{z}_{\backslash(j,n)},\boldsymbol{\beta})\triangleq\exp\Bigg{[}\sum_{m\in\mathcal{V}({n})}\beta^{j}_{xy}\,\delta\Big{(}{z}_{j,n}-{z}_{j,m}\Big{)}+$
$\sum_{p\in D(n)}\beta_{s}\,\delta\Big{(}{z}_{j,n}-{z}_{j+1,p}\Big{)}+\sum_{q\in U(n)}\beta_{s}\,\delta\Big{(}{z}_{j,n}-{z}_{j-1,q}\Big{)}\Bigg{]}.$
Using (9) and dropping all the terms independent of ${z}_{j,{n}}$,
@@
and the marginal density of $\ell_{j,n}$ is thus given by
@@
The multiscale labels are efficiently sampled using a checker-board parallelized strategy [41, 39], here adapted to work at multiple resolutions for the multiscale label prior, as illustrated in Fig.1-(b). The proposed strategy decomposes the set of labels at a given scale $j$ into two independent subsets, containing odd and even positions (represented in black and white in Fig.1-(b)). All labels within a subset are conditionally independent and can thus be sampled in parallel, with the labels of the other subset fixed, allowing odd and even labels to be sampled alternatively.
@@ 5.2 Sampling the Granularity Parameters $\beta_{s},\beta^{j}_{xy}$
It has been shown in [39] that it is possible to approximate the conditional density $f(\boldsymbol{\beta}|\mathbf{\Theta},\mathbf{x},\boldsymbol{\mu},\mathbf{z})=f(\boldsymbol{\beta}|\mathbf{z})$ by considering an auxiliary vector $\mathbf{w}$ that verifies $\mathbf{\Phi}(\mathbf{z})=\mathbf{\Phi}(\mathbf{w})$, where $\mathbf{\Phi}(\mathbf{z})$ is a tractable sufficient statistic. Then, by considering the Euclidean distance $\mathbf{\Phi}(\mathbf{z})-\mathbf{\Phi}(\mathbf{w})$, an approximation of the gradient can be computed. In practice, $\mathbf{w}$ is initialized with $\mathbf{z}$, and then updated with few Gibbs moves associated with the target density $f({\mathbf{z}}|\boldsymbol{\beta})$ (for more details see [40, 39]). This approximation of the gradient results in values around the neighborhood of $\boldsymbol{\beta}$ when choosing the correct statistic $\mathbf{\Phi}(\mathbf{z})$. Inspired by [39], we propose the statistic $\mathbf{\Phi}(\mathbf{z})=\sum_{j=j_{1}}^{j_{2}}\Phi_{\beta^{j}_{xy}}(\textbf{z}_{j})+\Phi_{\beta_{s}}(\textbf{z}_{j},\textbf{z}_{-j})$, for the scale granularity, where $\Phi_{\beta^{j}_{xy}}$ and $\Phi_{\beta_{s}}$ are defined in (25) and (26), and, for the spatial parameters, we use the statistic $\Phi_{\beta^{j}_{xy}}(\mathbf{z}_{j})$, $j=j_{1},...,j_{2}$. A truncation operation, denoted as $T_{[0,Q]}({x})$, is included to restrict the parameters to belong to the interval $(0,Q)$, $Q\in\mathbb{N}$. Following this heuristic but effective strategy, the granularity parameters are sampled as summarized in Algorithm 1. Intuitively, the sampler performs gradient descent to update the $\boldsymbol{\beta}$ parameters. It first generates auxiliary variables $\textbf{w}_{j}$ using Gibbs sampling. Then, the differences between the statistics evaluated in z and w are computed to approximate the intractable gradient, which are subsequently used to update the values of $\boldsymbol{\beta}$.
@@
Set $\mathbf{z}\leftarrow\mathbf{z}^{(t)}$, $[\beta_{s},\beta^{j_{1}}_{xy},...,\beta^{j_{2}}_{xy}]\leftarrow\boldsymbol{\beta}^{(t)}$
@@
       ${\textbf{w}}_{j}\leftarrow$ draw Gibbs moves of ${\textbf{z}}_{j}$ via (29), for $j=j_{1},...,j_{2}$ Set w=[$\textbf{w}=[\textbf{w}^{T}_{j_{1}},...,\textbf{w}^{T}_{j_{2}}]^{T}$ $\eta=10\,(t+r-1)^{-3/4}\big{(}\sum_{j=j_{1}}^{j_{2}}N_{j}\big{)}^{-1}$ Compute
@@
[…]
@@
The conditional distributions for the multifractal parameters contained in $\mathbf{\Theta}$ and the latent variables in ${\boldsymbol{\mu}}$ are:

### v2

@@ 3 Towards Multifractal-based Image Segmentation in the Fourier Domain
Model (10) provides an accurate estimation of $c_{2}$ (recall that $c_{2}={\theta}_{1}$). However, it is not designed for images with several multifractal textures. Here, we propose a new statistical model in the Fourier domain for modeling heterogeneous multifractal images. The limitation of (10) for computing $c_{2}$ is its reliance on a regular spatial grid. This prohibits its use for scenarios involving heterogeneous image regions with irregular shapes having different multifractal properties. We therefore develop a spectral likelihood that can handle multifractal regions with irregular shapes [fuentes2007approximate, Guillaumin2022], resulting in a more flexible model. It is important to note that this spectral likelihood is developed by means of the debiased spatial Whittle likelihood from [Guillaumin2022], used and adapted to address the complexity of computing the likelihood in the Fourier domain for irregularly shaped regions on regularly sampled images. Although the approach could be adapted to handle truly irregularly sampled data, this is beyond the scope of this work.
3.1 Likelihood on Regions With Irregular Shapes
Consider an image F that is composed of regions with different multifractal spectra. The log-leaders of F can be assumed to be a mixture of $K$ classes denoted as $\mathcal{C}_{k}$ with different multifractal parameters $\boldsymbol{\theta}_{k}\triangleq[{\theta}_{1}^{k},{\theta}_{2}^{k}]^{T}$, k=1,,Kk=1,...,K. For segmentation purposes, let $z_{j,n}$ be the label at position $n$ and scale $j$ that maps the observation ${\ell}_{j,n}$ to the class $\mathcal{C}_{k}$. The label vector at scale $j$ is denoted as $\textbf{z}_{j}=[z_{j,1},...,z_{j,N^{2}_{j}}]^{T}$ and ${\textbf{z}}=[\textbf{z}^{T}_{j_{1}},...,\textbf{z}^{T}_{j_{2}}]^{T}$. Let ${I}_{j,k}=\{{n}\in{N}^{2}_{j}\,|{z}_{j,n}=k\}$ be the subset of sites in $\textbf{z}_{j}$ with label $k$ at scale $j$, and $\mathcal{T}_{j,\mathbf{n},k}$ its corresponding indicator function:
@@ 3.1 Likelihood on Regions With Irregular Shapes
where $p(\ell_{j,\mathbf{n}\in I_{j,k}}|\boldsymbol{\theta}^{T}_{K})$ is the likelihood of a stationary zero mean Gaussian process with covariance (8) observed at locations $\mathbf{n}\in I_{j,k}$, and $\boldsymbol{\Theta}=[\boldsymbol{\theta}^{T}_{1},...,\boldsymbol{\theta}^{T}_{K}]^{T}$ contains the whole set of multifractal parameters. Just like (7), (15) is problematic for direct use in practice. In addition, obtaining an approximation similar to (10) for (15) is less straightforward.
@@ 3.1.1 Debiased Fourier Coefficients
Having different multifractality regions implies rewriting (10) using the coefficients that share the same multifractality, yielding an irregular shape which requires to adjust for bias in the Fourier transform [fuentes2007approximate, Guillaumin2022]. The log-leader Fourier coefficients are redefined for the $k$-th class label in terms of the irregular domain $\mathcal{T}_{j,{n},k}$ as
@@
where $\xi_{j,k}=1/\big({\sum_{\mathbf{n}}\mathcal{T}^{1/2}_{j,\mathbf{n},k}}\big)$ and $t_{j,k}=\xi^{-1}_{j,k}{\sum_{\mathbf{n}}\mathcal{T}_{j,\mathbf{n},k}\boldsymbol{\ell}_{j,\mathbf{n}}}$ are scaling constants considering the bias introduced by the irregular domain [fuentes2007approximate, Guillaumin2022].
@@ 3.1.2 Covariance Matrix
where $\mathbf{W}_{j,\textbf{n},k}$ is a weight matrix defined as [fuentes2007approximate, Guillaumin2022]:
@@
Accordingly, making use of data augmentation as in (10), the likelihood (20) can be written in a form that factorizes terms with the parameters ${\theta}^{k}_{1}$ and ${\theta}^{k}_{2}$, as
$p\big({\textbf{x}}_{k},{\boldsymbol{\mu}}_{k}|{\theta}^{k}_{1},{\theta}^{k}_{2},\textbf{z}\big)\propto\prod^{S}_{{s}=1}({\theta}^{k}_{2})^{-S}\exp\bigg(-\frac{1}{{\theta}^{k}_{2}}\boldsymbol{\mu}_{k}^{H}\mathbf{W}_{s,k}G^{-1}_{2,s}\boldsymbol{\mu}_{k}\bigg)({\theta}^{k}_{1})^{-S}\exp\bigg(-\frac{1}{{\theta}^{k}_{1}}({\textbf{x}}_{k}-{\boldsymbol{\mu}}_{k})^{H}\mathbf{W}_{s,k}G^{-1}_{1,s}({\textbf{x}}_{k}-\boldsymbol{\mu}_{k})\bigg),$ (21)
where $\mathbf{W}_{s,k}\triangleq\mathbf{W}_{j,\textbf{n},k}$, and ${\boldsymbol{\mu}}_{k}=[{{\mu}}_{1,k}^{T},...,{{\mu}}_{S,k}^{T}]^{T}$ is the complex-valued latent vector. The likelihood for all classes is finally defined as:
@@ 4 Bayesian Model for Multifractal Segmentation
With the likelihood (22), the multifractal parameters and labels corresponding to image regions with different regularities can be estimated using Bayesian estimators. This approach uses Bayesian inference, where prior distributions are assigned to the unknown model parameters, and these parameters are estimated using the resulting posterior distribution. In this section, the proposed Bayesian model for image segmentation and multifractal parameter estimation is presented.
@@ 4.0.1 Multiscale Label Prior
To account for the multiscale nature of the log-leaders, the label dependencies within and across scales are modeled via a graph. Specifically, inter-scale connections link neighbors across scales, and the intra-scale connections model the spatial correlation of labels. Figure 2-(a) illustrates the proposed graph structure at a given scale $j$, where the labels to be sampled at position $n$ (white circle) use the values of its neighbors in space and scale (gray dots) to update their values accordingly. Formally, the corresponding multiscale prior for $\textbf{z}_{j}\in\{1,...,K\}^{n}$ at scale $j$ is defined as a modified version of the classical Potts-Markov field [wu1982potts],
@@
The potential for the spatial (intra-scale) dependencies is:
@@
where $\mathcal{V}(n)$ contains the four spatial neighbors of $n$ [pereyra2012segmentation], as depicted in Fig. 2-(a). The potential for the scale (inter-scale) dependencies is defined as:
$\Phi_{\beta_{s}}(\textbf{z}_{j},\textbf{z}_{-j})\triangleq\sum_{n=1}^{N^{2}_{j}}\sum_{p\in D(n)}\beta_{s}\,\delta\Big({z}_{j,n}-\textbf{z}_{j+1,p}\Big)+\sum_{q\in U(n)}\beta_{s}\,\delta\Big({z}_{j,n}-\textbf{z}_{j-1,q}\Big),$ (26)
where $D(n)$ and $U(n)$ are multiscale index mapping functions that map the index $n$ to its corresponding parent (down) or child (up) coefficient at a given scale, respectively (see Fig. 2-(a)). The inter-scale and spatial granularity coefficients $\beta_{s}$ and $\beta^{j}_{xy}$, respectively, adjust the degree of homogeneity across scales and in space, and are gathered in the vector $\boldsymbol{\beta}=[\beta_{s},\beta^{1}_{xy},...,\beta^{J}_{xy}]$. Finally, $\tilde{C}=C(\beta_{s},\beta^{j}_{xy})$ is a partition function defined as
@@
which is generally intractable because the number of summation terms grows exponentially with the size of $\textbf{z}_{j}$ [pereyra2013estimating].
@@ 4.1.1 Granularity Parameters $\beta_{s}$ and $\beta^{j}_{xy}$
Fixing a specific value for the parameters $\beta_{s}$ and $\beta^{j}_{xy}$ in (24) is a difficult task due to individual image’s spatial structure and organization that can change from one image to another. Indeed, small values of $\beta^{j}_{xy}$ lead to noisy segmentation whereas large ones lead to coarse or over-smoothed regions. Inspired by [pereyra2014maximum, pereyra2013estimating], the granularity parameters are jointly estimated with the labels and are assigned a uniform prior distribution on the interval $(0,Q)$, i.e., $f(\boldsymbol{\beta})=\mathcal{U}_{(0,Q)}(\boldsymbol{\beta})$, where $\boldsymbol{\beta}=[\beta_{s},\beta^{1}_{xy},...,\beta^{J}_{xy}]$ and $Q$ is a given constant related to the number of classes $K$.
4.1.2 Multifractal Parameter Prior
The natural conjugate priors for the parameters $\boldsymbol{\theta}^{k}_{i}$ are inverse-gamma distributions $\mathcal{IG}(\alpha_{i,k},\gamma_{i,k})$ [wendt2018multifractal, combrexelle2015bayesian], with fixed constants $\alpha_{i,k},\gamma_{i,k}$ $(i=1,2)$ defined, here, for each multifractal region k=1,,Kk=1,...,K.
4.2 Posterior Distribution
@@
The posterior distribution of the model parameters depends on the spectral likelihood (22) and the priors as follows:
@@
The Bayesian model is summarized in Fig. 3. Note that given the complexity of the posterior distribution (28), closed-form expressions for the estimators of the unknown parameters $\boldsymbol{\Theta}$, z and $\boldsymbol{\beta}$ cannot be obtained. As an alternative, it is possible to generate samples that are asymptotically distributed according to the target distribution (28) [pereyra2013estimating]. These samples are then used to approximate Bayesian estimators of the unknown model parameters.
@@ 5 Sampling Procedure
The vector z cannot be sampled jointly since its entries are correlated by the Potts prior. The conditional distribution of a label $z_{j,n}$, $1<j<J$, given all the other labels $\textbf{z}_{\backslash(j,n)}$, is
${z}_{j,{n}}\,|\,\textbf{z}_{\backslash(j,n)},{\textbf{x}},{\boldsymbol{\mu}},\boldsymbol{\Theta},\boldsymbol{\beta}\propto\!p({z}_{j,{n}}|\textbf{z}_{\backslash(j,n)},\boldsymbol{\beta})\,p\big({\textbf{x}},{\boldsymbol{\mu}}|\boldsymbol{\Theta},\textbf{z}_{\backslash(j,n)}\big)$ (29)
@@
$p({z}_{j,{n}}|\textbf{z}_{\backslash(j,n)},\boldsymbol{\beta})\triangleq\exp\Bigg[\sum_{m\in\mathcal{V}({n})}\beta^{j}_{xy}\,\delta\Big({z}_{j,n}-{z}_{j,m}\Big)+\sum_{p\in D(n)}\beta_{s}\,\delta\Big({z}_{j,n}-{z}_{j+1,p}\Big)+\sum_{q\in U(n)}\beta_{s}\,\delta\Big({z}_{j,n}-{z}_{j-1,q}\Big)\Bigg].$
Using (9) and dropping all the terms independent of ${z}_{j,{n}}$, the following result is obtained:
@@
and the marginal density of $\ell_{j,n}$ is:
@@
The multiscale labels are efficiently sampled using a checker-board parallelized strategy [gonzalez2011parallel, pereyra2013estimating], here adapted to work at multiple resolutions for the multiscale label prior, as illustrated in Fig.2-(b). The proposed strategy decomposes the set of labels at a given scale $j$ into two independent subsets, containing odd and even positions (represented in black and white in Fig.2-(b)). All labels within a subset are conditionally independent and can thus be sampled in parallel, with the labels of the other subset fixed, allowing odd and even labels to be sampled alternatively.
@@ 5.2 Sampling the Granularity Parameters $\beta_{s},\beta^{j}_{xy}$
It has been shown in [pereyra2013estimating] that it is possible to approximate the conditional density $f(\boldsymbol{\beta}|\mathbf{\Theta},\mathbf{x},\boldsymbol{\mu},\mathbf{z})=f(\boldsymbol{\beta}|\mathbf{z})$ by considering an auxiliary vector $\mathbf{w}$ that verifies $\mathbf{\Phi}(\mathbf{z})=\mathbf{\Phi}(\mathbf{w})$, where $\mathbf{\Phi}(\mathbf{z})$ is a tractable sufficient statistic. Then, by considering the Euclidean distance $\mathbf{\Phi}(\mathbf{z})-\mathbf{\Phi}(\mathbf{w})$, an approximation of the gradient can be computed. In practice, $\mathbf{w}$ is initialized with $\mathbf{z}$, and then updated with few Gibbs moves associated with the target density $f({\mathbf{z}}|\boldsymbol{\beta})$ (for more details see [pereyra2014maximum, pereyra2013estimating]). This approximation of the gradient results in values around the neighborhood of $\boldsymbol{\beta}$ when choosing the correct statistic $\mathbf{\Phi}(\mathbf{z})$. Inspired by [pereyra2013estimating], we propose the statistic $\mathbf{\Phi}(\mathbf{z})=\sum_{j=j_{1}}^{j_{2}}\Phi_{\beta^{j}_{xy}}(\textbf{z}_{j})+\Phi_{\beta_{s}}(\textbf{z}_{j},\textbf{z}_{-j})$, for the scale granularity, where $\Phi_{\beta^{j}_{xy}}$ and $\Phi_{\beta_{s}}$ are defined in (25) and (26), and, for the spatial parameters, we use the statistic $\Phi_{\beta^{j}_{xy}}(\mathbf{z}_{j})$, $j=j_{1},...,j_{2}$. A truncation operation $T_{[0,Q]}({x})$ is introduced to restrict the parameters to belong to the interval $(0,Q)$, $Q\in\mathbb{N}$. Following this heuristic but effective strategy, the granularity parameters are sampled as summarized in Algorithm 1. Intuitively, the sampler performs gradient descent to update the $\boldsymbol{\beta}$ parameters. It first generates auxiliary variables $\textbf{w}_{j}$ using Gibbs sampling. Then, the differences between the statistics evaluated in z and w are computed to approximate the intractable gradient, which are used to update the values of $\boldsymbol{\beta}$.
The conditional distributions of the multifractal parameters contained in $\mathbf{\Theta}$ and the latent variables in $\boldsymbol{\mu}$} are:
@@
Set $\mathbf{z}\leftarrow\mathbf{z}^{(t)}$, $\beta_{s},\beta^{j_{1}}_{xy},...,\beta^{j_{2}}_{xy}]\leftarrow\boldsymbol{\beta}^{(t)}$
@@
  ${\textbf{w}}_{j}\leftarrow$ draw Gibbs moves of ${\textbf{z}}_{j}$ via (29), for $j=j_{1},...,j_{2}$ Set $\textbf{w}=[\textbf{w}^{T}_{j_{1}},...,\textbf{w}^{T}_{j_{2}}]^{T}$ $\eta=10\,(t+r-1)^{-3/4}\big(\sum_{j=j_{1}}^{j_{2}}N_{j}\big)^{-1}$ Compute
@@
return $\boldsymbol{\beta}=[\beta_{s},\beta^{j_{1}}_{xy},...,\beta^{j_{2}}_{xy}]$